OpenStudy (anonymous):

Prove sinx^2= 1-cos2x/2, use the left side to start the proof.....help please?

OpenStudy (mertsj):

\[\sin x^2=1-\frac{\cos (2x)}{2}\]

OpenStudy (mertsj):

Is that what you meant to post?

OpenStudy (anonymous):

wait instead of sinx^2 its actually sin^2x sorry my bad

OpenStudy (mertsj):

\[\sin ^2x=1-\frac{\cos 2x}{2}\]

OpenStudy (mertsj):

Like that?

OpenStudy (anonymous):

it looks like this: